x^2+y^2+6x-8y-155=0
Find the radius and the center of the circle. For the radius use radicals as needed for an exact answer. For the center type in an ordered pair using fractions and integers as needed.

Answer by lwsshak3(11628)

Find the radius and the center of the circle.
Equation of a circle:(x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
x^2+y^2+6x-8y-155=0
complete the squares
x^2+6x+y^2-8y-155=0
(x^2+6x+9)+(y^2-8y+16)=155+9+16
(x+3)^2+(y-4)^2=180
center: (-3,4)
radius √180